autumnspivey:

Solve for x: one over four (4x + 15) = 24 (5 points) 49 two hundred seven over four eighty one over four 9

umm:

Given\[\frac{ 1 }{ 4 }(4x+15)=24\]Simplify both sides of the equation and distribute \[\frac{ 1 }{ 4 }(4x+15)=24\]\[\left(\begin{matrix}\frac{ 1 }{ 4 }\end{matrix}\right)(4x)+\left(\begin{matrix}\frac{ 1 }{ 4 }\end{matrix}\right)(15)=24\]\[x+\frac{ 15 }{ 4 }=24\]Subtract ¹⁵⁄₄ from both sides \[x+\frac{ 15 }{ 4 }-\frac{ 15 }{ 4 }=24-\frac{ 15 }{ 4 }\]\[x=\frac{ 81 }{ 4 }\]Therefore, your answer is C\[x=\frac{ 81 }{ 4 }\]

umm:

Also, welcome to QuestionCove, @autumnspivey.

darkknight:

Also another way to do this \[1/4(4x+15)=24\] Multiply both sides by 4 \[4x+15=96\] \[4x+15-15=96-15\] \[4x=81\] \[x=81/4\]

bemine11:

2.25 i think